A solution contains 10.0 g of carbon tetrachloride, CCl4, and 20.0 g of benzene, C6H6. Calculate the mole fraction of each component.
1
Expert's answer
2020-12-14T14:34:50-0500

MM(CCl4) = 153.82 g/mol

n(CCl4) "= \frac{10.0}{153.82} = 0.065 \;mol \\"

MM(C6H6) = 78.11 g/mol

n(C6H6) "= \frac{20.0}{78.11} = 0.256 \;mol"

Total moles = 0.065 + 0.256 = 0.321 mol

Proportion:

0.321 – 1

0.065 – x

"x = \frac{0.065 \times 1}{0.321} = 0.2025"

The mole fraction of CCl4 is 0.2025

The mole fraction of C6H6 is (1 – 0.2025) = 0.7975